OpenStudy (lordhelix8th):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {43, 42, 56, 62, 72} Data Set 2: {76, 57, 81, 51, 70}

OpenStudy (lordhelix8th):

@rational @EclipsedStar

OpenStudy (lordhelix8th):

@chosenmatt

OpenStudy (lordhelix8th):

@Loser66

OpenStudy (lordhelix8th):

@Imperfect_Queen @Teddyiswatshecallsme

OpenStudy (lordhelix8th):

@TheAsker2002

OpenStudy (lordhelix8th):

@tkhunny

OpenStudy (anonymous):

2.4

OpenStudy (loser66):

set 1: mean: 55 , MAD : 10

OpenStudy (loser66):

set 2 : mean: 67, MAD : 10.4

OpenStudy (loser66):

Woah!!! the concept is totally new to me. Thanks for asking. It enlarges my knowledge to the field I never known before.

OpenStudy (lordhelix8th):

yw

OpenStudy (lordhelix8th):

lol

OpenStudy (loser66):

Method: set 1 : find the mean out by (43+42+56+62+72)/5 = 55 find MAD by formula: \(\dfrac{1}{n}\sum_{i=1}^n |x_i-\bar x|\), where \(\bar x\) is the mean above, that is 55 Hence MAD=\(\dfrac{1}{5} [|43-55|+|42-55|+56-55|+|62-55|+|72-55|]=10\) lalalalalala.... I know something new,

OpenStudy (loser66):

Do the same with set 2.

OpenStudy (lordhelix8th):

so the answer is 10

OpenStudy (loser66):

I gave you both of them above the method 1) 10 2) 10.4

OpenStudy (lordhelix8th):

so the answer is 10.4

OpenStudy (lordhelix8th):

thank you